# CAPE: Covariate-Adjusted Pre-Training for Epidemic Time Series Forecasting

### v2

@@ 3.1.2 Model Instantiation
Latent Environment Estimator $q_{\phi}(Z|X)$. We model $p(Z|X)$ using a latent environment estimator $q_{\phi}(Z|X)$. Since environmental influences vary over time, we apply patching Nie et al. (2022) to manage granularity in environment estimation. This prevents overly specific or generalized estimations that could obscure key temporal fluctuations. We divide the input $\mathbf{x}$ into $C$ non-overlapping patches, $\mathbf{x}=[\mathbf{x}_{1},\dots,\mathbf{x}_{C}]$, where $\mathbf{x}_{c}\in\mathbb{R}^{T/C}$. Then, a self-attention layer $f_{\text{enc}}$ captures temporal dependencies between patches, producing contextualized representations $\mathbf{h}^{(l)}_{c}=f_{\text{enc}}(\mathbf{x}^{(l)}_{c})$ for each patch at layer $l$. Subsequently, since the environment influences only the spurious component of the input, we introduce a transformation $\mathbf{W}_{s}^{(l)}$ to capture the spurious component of $\mathbf{h}^{(l)}_{c}$. Finally, we model $q_{\phi}(Z|X)$ as a cross-attention layer that captures the relation between each patch and the latent environment representations:
@@
where $\pi_{k,c}^{(l)}$ is the output probability of the environment $z_{k}$ for the $c$-th patch, and $\mathbf{W}_{k}^{(l)}$ is a transformation layer for $\mathbf{e}_{k}$. Such operation not only takes into account the contextualized representation of the current time period, but also considers the latent environment representations, which made it possible to infer the densities of other environment distributions with different latent representations.
Epidemic Predictor $p_{\Theta}(\hat{Y}|X,Z)$. Unlike previous studies, which do not explicitly model environment states, we incorporate these states directly into the input using their latent representations $\mathbf{e}_{k}$. Specifically, we model the predictor $p_{\Theta}(\hat{Y}|X,Z)$ by employing a weighted sum over the combined representations of each environment and the input using Hadamard product, i.e., $f_{\text{enc}}(\mathbf{x}_{c}^{(l)})\odot\mathbf{e}_{k}$. Finally, we apply a feed-forward layer to compute the output representations, serving as the input for the next layer. Integrating these components, the CAPE encoder can be expressed as:
$\small\mathbf{x}^{(l+1)}_{c}=\sigma\left(\mathbf{W}_{f}^{(l)}\sum_{k=1}^{K}\pi_{k,c}^{(l)}\left[f_{enc}({\mathbf{x}_{c}^{(l)}})\odot\mathbf{e}_{k}\right]\right),$ (6)
@@ 3.2 Pre-training Objectives for Epidemic Forecasting
Random Masking with Environment Estimation. To capture features from large unlabeled epidemic time series data, we employ a masked time series modeling task Kamarthi & Prakash (2023); Goswami et al. (2024) (Figure 2(c)) that masks 30% of input patches. As depicted in Figure 2, the generation of $X$ depends on the environment $Z$, indicating that accurate patch reconstruction requires capturing both temporal and environmental dependencies. Unlike prior studies that overlook the environment’s role, we utilize an environment estimator $q_{\phi}(Z|X)$ to infer $Z$, aiding both reconstruction and estimator training. During pre-training, input $\mathbf{x}$ is transformed into masked input and label pairs $(\tilde{\mathbf{x}},\mathbf{x})$, with the original time series serving as label $y$. The reconstruction $\hat{\mathbf{x}}=h_{\psi}(g_{\theta}(\tilde{\mathbf{x}}))$ is optimized using Mean Squared Error (MSE): $\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})=\text{MSE}(\hat{\mathbf{x}},\mathbf{x})$.
Hierarchical Environment Contrasting. Two consecutive time series samples, ${\bf x}$ and ${\bf x}^{\prime}$, can include overlapping regions when divided into multiple patches. These overlapping patches, although identical, can exhibit contextual variations influenced by their different adjacent patches. As indicated by Eq. (5), such variations can alter the latent patch-wise representations, leading to inconsistencies in the environmental estimates for the same patch across the samples. To ensure that each patch’s environment remains context-invariant, we propose a hierarchical environment contrasting scheme inspired by Yue et al. (2022). We define an aggregated latent environment representation $\hat{\mathbf{e}}^{(l)}_{c}=\sum_{k=1}^{K}\mathbf{e}_{k}\pi_{k,c}^{(l)}$ to represent the weighted environment states for the $c$-th patch. For contrastive loss computation, we use the combined representation $\hat{\mathbf{E}}_{j,c}^{(l)}=\sigma(\mathbf{W}_{f}^{(l)}(\hat{\mathbf{e}}^{(l)}_{c}\odot\mathbf{h}^{(l)}_{c}))$ for $c$-th patch of sample $j$. Additionally, $\hat{\mathbf{E}^{\prime}}_{j,c}^{(l)}$ denotes the representation in the context of $\mathbf{x}^{\prime}$. Finally, we compute a patch-wise contrastive loss:
@@
$+\log\left(\sum_{b\in B}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}^{\prime}}_{(b,c)}\right)+\mathbb{I}_{j\neq b}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}}_{(b,c)}\right)\right)$
$+\log\left(\sum_{t\in\Omega}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}^{\prime}}_{(j,t)}\right)+\mathbb{I}_{c\neq t}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}}_{(j,t)}\right)\right).$
where $B$ is the batch size, $\Omega$ denotes the overlapping patches, and $\mathbb{I}$ is the indicator function. The above equation contains three key terms: (1) The first term encourages the representations of the same patch from two different contexts to be similar, which preserves the context-invariant nature of environments. (2) The second term (Instance-wise Contrasting) treats $\hat{\mathbf{e}}^{(l)}_{c}$ from different samples in the batch as negative pairs, which promotes dissimilar representations, and enhances diversity among instances. (3) The third term (Temporal Contrasting) treats the representations of different patches from overlapping regions ($\Omega$) as negative pairs, which encourages differences across temporal contexts.
Pre-Training Loss. Given a batch of $B$ samples $\mathbf{X}\in\mathbb{R}^{B\times T}$, we combine the reconstruction loss and the contrastive loss, yielding the final loss function for pre-training:
$\mathcal{L}_{\text{final}}=\sum\nolimits_{\mathbf{x}\in\mathbf{X}}\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})+\alpha\,\mathcal{L}_{\text{CL}}(\hat{\mathbf{E}}^{(L)},\hat{\mathbf{E}}^{\prime(L)}),\;\mathbf{X}\sim\mathcal{D}_{\text{pre}}$
where $L$ is the number of layers, and $\alpha$ is the hyperparameter used to balance the contrastive loss and the reconstruction loss. Further analysis can be found in Appendix A.10.
@@ 3.3 Optimization of the CAPE Framework
To effectively maximize the variational lower bound in Eq. (4), we employ the Expectation-Maximization (EM) algorithm to iteratively update the latent environments and epidemic predictor. The pseudo algorithm for the optimization procedure is provided in Appendix A.3.
E-Step: Estimating Latent Environments. In the E-step, we aim to identify the environment states $Z$ and the corresponding distribution $p(Z)$ that result in the target distribution $p(Y)$. This involves maximizing the expected likelihood of $p(Y|Z)$ given $p(Z)$. We freeze the epidemic predictor $p_{\Theta}(\hat{Y}|X,Z)$ and the environment estimator $q_{\phi}(Z|X)$, treating them as oracles, which means $p_{\Theta}(\hat{Y}|X,Z)=p(Y|X,Z)$ and $q_{\phi}(Z|X)=q(Z|X)$. While actively updating the environment representations $\mathbf{E}=[\mathbf{e}_{1},\mathbf{e}_{2},...\mathbf{e}_{k}]$, the optimization of the environment states $Z$ is learned through maximizing $\mathbb{E}_{p(Z)}[p(Y|Z)]=\mathbb{E}_{p(X)}[\mathbb{E}_{q_{\phi}(Z|X)}p_{\Theta}(Y|X,Z)]$, which is equivalent to minimizing the expected reconstruction loss:
$\mathbf{E}^{t+1}\leftarrow\arg\min\nolimits_{\mathbf{E}}\left[\mathbb{E}_{\mathbf{x}\sim p(X)}[\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})]\right].$ (7)
We use subscript $t$ to denote the pre-update distribution and derive the updated distribution $p^{t+1}(Z)$ as $q^{t+1}_{\phi_{t}}(Z)$, along with the updated environment representations $\mathbf{E}^{t+1}$.
M-Step: Optimizing Epidemic Predictor. In the M-step, we aim to optimize the epidemic predictor by maximizing its predictive power and regularizing the environment distribution. During this step, the environment representations $\mathbf{E}^{t+1}$ are held fixed. We have the following theorem:
@@
The detailed proof can be found in Appendix A.1. Theorem 3.2 indicates that the optimization of the model’s predictive ability can be approximated by Eq. (8), which corresponds to the expectation of $\mathcal{L}_{recon}$. To further enhance robustness, the contrastive loss is combined to regularize the environment estimator. Therefore, the overall optimization objective becomes minimizing the final pre-training loss:
$\Theta_{t+1}\leftarrow\arg\min\nolimits_{\Theta}\left[\mathcal{L}_{\text{final}}(\mathbf{X},\hat{\mathbf{X}},\mathbf{E}^{t+1})\right].$ (9)
@@
Table 2: Few-shot learning results with horizons ranging from 1 to 16 future steps. The length of the lookback window is set to 36. Each model is evaluated after being trained on 20%, 40%, 60%, and 80% of the full training data. $\Delta(\%)$ stands for the relative improvement of the model after training with 20% more data in terms of average MSE over all horizons. The full result is shown in Appendix A.5.
Dataset/Model CAPE PatchTST Dlinear MOMENT PEM
@@ 4.1 Experiment Setup
Datasets. For pre-training CAPE, PatchTST, and PEM, we manually collected 17 distinct weekly-sampled diseases from Project Tycho van Panhuis et al. (2018). For evaluation, we utilize five downstream datasets covering various diseases and locations: ILI USA Centers for Disease Control and Prevention (2023a), ILI Japan National Institute of Infectious Diseases (2023), COVID-19 USA Dong et al. (2020), Measles England Lau et al. (2020a), and Dengue across countries OpenDengue (2023). Additionally, RSV Centers for Disease Control and Prevention (2023c) and Monkey Pox Centers for Disease Control and Prevention (2023b) infections in the US are used to test zero-shot performance. More details can be found in Appendix A.2.
Baselines. For baselines, we leverage the models from the comprehensive toolkit EpiLearn Liu et al. (2024a). To provide a comprehensive evaluation, we compare CAPE with two sets of models: non-pretrained and pre-trained. Non-pretrained models include statistical methods like ARIMA Panagopoulos et al. (2021), RNN-based Wang et al. (2020); Natarajan et al. (2023) approaches such as LSTM and GRU, the linear model DLinear Zeng et al. (2023), and transformer-based methods Wu et al. (2021); Zhou et al. (2021, 2022). For pre-trained models, we evaluate popular approaches including PatchTST Nie et al. (2022), PEM Kamarthi & Prakash (2023), and a time series foundation model MOMENT Goswami et al. (2024). More experimental details can be found in Appendix A.3.
4.2 Baseline Comparison
We now evaluate the CAPE model under three settings: fine-tuning, few-shot fine-tuning, and zero-shot forecasting.
4.2.1 Fine-Tuning (Full-Shot Setting)
@@
4.2.2 Few-Shot and Zero-Shot Performance
@@
Table 3: Zero-shot performance with a lookback window length of 12. All results are averaged over 4 weeks or days in the future. $\Delta(\%)$ stands for the relative improvement of CAPE over the baselines.
@@
4.3 Ablation Study
We conducted an ablation study to assess CAPE’s components (Table 4). Replacing environment estimators with non-disentangling self-attention layers consistently worsened performance across all datasets, notably increasing ILI USA’s MSE from 0.309 to 0.532, underscoring the importance of environmental factors. Similarly, removing contrastive loss while retaining environment estimators raised Measles’ MSE from 0.269 to 0.344, with smaller increases for ILI USA and Dengue. Training CAPE directly on downstream datasets without pre-training also decreased performance, with MSE rising to 0.319 (ILI USA), 0.326 (Measles), and 1.162 (Dengue), though less than removing environment estimation. These results indicate that all CAPE components are essential for optimal forecasting and that tailoring component emphasis to dataset characteristics can further enhance performance.
Table 4: Ablation study of removing components from CAPE.

### v3

@@ 3.2 Model Instantiation
Latent Environment Estimator $q_{\phi}(Z|X)$. We approximate the distribution of environmental contexts $p(Z)$ using a latent environment estimator $q_{\phi}(Z|X)$. Epidemic infection trajectories are profoundly shaped by dynamic environmental influences—such as the implementation and lifting of non-pharmaceutical interventions (NPIs), seasonal drivers affecting pathogen viability (e.g., influenza seasonality), or shifts in population behavior—which can vary significantly over different periods of an outbreak. Modeling the impact of these potentially complex and time-varying influences across an entire long infection trajectory can be inefficient and may obscure shorter-term but critical environmental shifts. Therefore, we apply patching (Nie et al., 2022) to manage the granularity in our environment estimation. This approach allows the model to identify distinct environmental ”regimes” or periods within an epidemic wave that might correspond to different dominant external factors (e.g., a period of strict lockdown versus a period of reopening). Such granularity helps prevent overly coarse or overly fine-grained estimations that could obscure epidemiologically relevant shifts in transmission dynamics, such as changes in the effective reproduction number or the shape of the epidemic curve. We divide the historical infection input $\mathbf{x}$ (e.g., weekly case counts) into $C$ non-overlapping temporal patches, $\mathbf{x}=[\mathbf{x}_{1},\dots,\mathbf{x}_{C}]$, where $\mathbf{x}_{c}\in\mathbb{R}^{T/C}$. A self-attention layer $f_{\text{enc}}$ then captures temporal dependencies between these patches, producing contextualized representations of the local infection trend $\mathbf{h}^{(l)}_{c}=f_{\text{enc}}(\mathbf{x}^{(l)}_{c})$ for each patch at layer $l$. Subsequently, the environment state and the patch representation are projected to the same dimension and we calculate the cross-attention score using a softmax function as the environment probability:
@@
where $\pi_{k,c}^{(l)}=q_{\phi}(Z=z_{k}|X=x_{c})$ is the output probability of the environment $z_{k}$ for the $c$-th patch, and $\mathbf{W}_{k}^{(l)}$ and $\mathbf{W}_{s}^{(l)}$ are learnable linear layers. This mechanism allows the model to not only leverage the observed infection pattern within the patch $\mathbf{h}_{c}^{(l)}$ but also to consider a set of learned environmental influences $\textbf{e}_{k}$. The $\textbf{e}_{k}$ are optimized to capture distinct types of environmental conditions, expressed as a vector of channel-wise weights, relevant to disease spread across many historical epidemics encountered during pre-training. This design is crucial as it enables the model to potentially infer the influence of combinations or intensities of environmental factors not explicitly seen together in the training data for a specific downstream task, thereby enhancing its ability to generalize to novel epidemic scenarios or evolving outbreaks.
Epidemic Predictor $p_{\Theta}(\hat{Y}|X,Z)$. Unlike previous studies, which do not explicitly model environment states, we incorporate these states directly into the patched infection trend using their latent representations $\mathbf{e}_{k}$. Specifically, we model the predictor $p_{\Theta}(\hat{Y}|X,Z)$ in Eq (2) by employing a weighted sum over the combined representations of each environment and the infection trend using Hadamard product, i.e., $f_{\text{enc}}(\mathbf{x}_{c}^{(l)})\odot\mathbf{e}_{k}$. Finally, we apply a feed-forward layer is applied. This step gives us the adjusted environment-invariant representation for $X_{ca}$ in Figure 2, and we conduct such adjustment in a layer-wise manner, which can be expressed as:
$\small\mathbf{x}^{(l+1)}_{c}=\sigma\left(\mathbf{W}_{f}^{(l)}\sum\nolimits_{k=1}^{K}\pi_{k,c}^{(l)}\left[f_{enc}({\mathbf{x}_{c}^{(l)}})\odot\mathbf{e}_{k}\right]\right),$ (5)
@@ 3.3 Pre-training Objectives for Epidemic Forecasting
Random Masking with Environment Estimation. To learn from large unlabeled epidemic time series, we adopt a masked modeling task (Kamarthi & Prakash, 2023; Goswami et al., 2024), masking 30% of input patches (Figure 3(c)). As shown in Figure 2, input $X$ is influenced by the environment $Z$, so reconstructing patches requires modeling both temporal and environmental dependencies. Unlike previous methods that ignore this, we use backdoor adjustment via an environment estimator $q_{\phi}(Z|X)$ to obtain environment-invariant representations. During pre-training, the input $\mathbf{x}$ is masked to produce pairs $(\tilde{\mathbf{x}},\mathbf{x})$, with the reconstruction $\hat{\mathbf{x}}=h_{\psi}(g_{\theta}(\tilde{\mathbf{x}}))$ trained using Mean Squared Error: $\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})=\text{MSE}(\hat{\mathbf{x}},\mathbf{x})$.
Hierarchical Environment Contrasting for Robust Epidemiological Contextualization. During optimization, inconsistent environment representations can occur if the model, when processing two infection trajectories (${\bf x}$ and ${\bf x}^{\prime}$) that share an identical overlapping time period, estimates different environmental contexts for that common period based on the distinct temporal context. To address the challenge and improve consistency in representing environments for adjacent epidemic time series segments, we propose Hierarchical Environment Contrasting. Specifically, We define an aggregated environment representation as the confluence of environmental factors, $\hat{\mathbf{e}}^{(l)}_{c}=\sum_{k=1}^{K}\mathbf{e}_{k}\pi_{k,c}^{(l)}$, for each patch $c$, allowing it to reflect a nuanced, weighted combination of $K$ learned latent environments ($\mathbf{e}_{k}$) rather than a single discrete state. Our contrastive feature, $\hat{\mathbf{E}}_{j,c}^{(l)}=\sigma(\mathbf{W}_{f}^{(l)}(\hat{\mathbf{e}}^{(l)}_{c}\odot\mathbf{h}^{(l)}_{c}))$, is designed to capture this essential interaction. It explicitly combines the aggregated environment representation ($\hat{\mathbf{e}}^{(l)}_{c}$) with the local infection trend representation ($\mathbf{h}^{(l)}_{c}$) via a Hadamard product ($\odot$), enabling the model to learn how specific environmental profiles alter observed infection trends (e.g., by changing growth rates). The overarching goal, realized through the contrastive loss (Equation 3.3), is to ensure these learned features $\hat{\mathbf{E}}_{j,c}^{(l)}$ are consistent across genuinely comparable epidemiological contexts and distinct for dissimilar underlying situations, even if superficial case counts appear similar, fostering robust, discriminative representations of environmental influence. The patch-wise contrastive loss is:
@@
$\hskip 28.45274pt+\log\left(\sum\nolimits_{b\in B}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}^{\prime}}_{(b,c)}\right)+\mathbb{I}_{j\neq b}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}}_{(b,c)}\right)\right)$
$\hskip 28.45274pt+\log\left(\sum\nolimits_{t\in\Omega}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}^{\prime}}_{(j,t)}\right)+\mathbb{I}_{c\neq t}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}}_{(j,t)}\right)\right).$
where $B$ is the batch, $\Omega$ is the set of overlapping patches, and $\mathbb{I}$ is the indicator function. For more details, please refer to Appendix A.4.
Given a batch $\mathbf{X}\in\mathbb{R}^{B\times T}$, our pre-training loss is: $\mathcal{L}_{\text{final}}=\sum\nolimits_{\mathbf{x}\in\mathbf{X}}\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})+\alpha\,\mathcal{L}_{\text{CL}}(\hat{\mathbf{E}}^{(L)},\hat{\mathbf{E}}^{\prime(L)})$, where $L$ is the number of layers and $\alpha$ balances reconstruction and contrastive losses.
@@ 3.4 Optimization of the CAPE Framework
To maximize the variational lower bound in Eq. (3), we adopt the Expectation-Maximization (EM) algorithm to iteratively update latent environments and the epidemic predictor (see Appendix A.5 for the pseudo algorithm).
E-Step: Estimating Latent Environments. We estimate the environment states $Z$ and their distribution $p(Z)$ by maximizing the expected likelihood of $p(Y|Z)$. Fixing the epidemic predictor $p_{\Theta}(\hat{Y}|X,Z)$ and environment estimator $q_{\phi}(Z|X)$, we treat them as oracles such that $p_{\Theta}(\hat{Y}|X,Z)=p(Y|X,Z)$ and $q_{\phi}(Z|X)=p(Z)$. The environment representations $\mathbf{E}$ are updated by minimizing the expected reconstruction loss: $\mathbf{E}^{t+1}\leftarrow\arg\min\nolimits_{\mathbf{E}}\left[\mathbb{E}_{\mathbf{x}\sim p(X)}\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})\right]$. This yields the updated distribution $p^{t+1}(Z)=q^{t+1}_{\phi_{t}}(Z)$ and representations $\mathbf{E}^{t+1}$.
M-Step: Optimizing Epidemic Predictor. With $\mathbf{E}^{t+1}$ fixed, we provide the following theorem:
@@
Assuming $q^{t+1}_{\phi_{t}}(Z)=p^{t+1}(Z)$ and an L2 norm is applied on $\phi$, the variational lower bound in Eq. (3) can be approximated as follows: $\mathbb{E}_{p(X)}\left[\mathbb{E}_{q^{t+1}_{\phi_{t}}(Z|X)}\left[\log p^{t+1}_{\Theta_{t+1}}(\hat{Y}|X,Z)\right]\right]-C$, which is equivalent to minimizing the expected reconstruction loss $\mathbb{E}_{\mathbf{x}\sim p(X)}[\mathcal{L}_{recon}(\mathbf{x},\hat{\mathbf{x}})]$
Further theoretical details are provided in Appendix A.3. Therefore, we optimize model parameters by minimizing the expected reconstruction loss and incorporating a contrastive loss to regularize the environment estimator: $\Theta_{t+1}\leftarrow\arg\min\nolimits_{\Theta}\left[\mathcal{L}_{\text{final}}(\mathbf{X},\hat{\mathbf{X}},\mathbf{E}^{t+1})\right]$.
@@
Table 2: Few-shot learning results with horizons ranging from 1 to 16 future steps. The length of the lookback window is set to 36. Each model is evaluated after being trained on 20%, 40%, 60%, and 80% of the full training data. $\Delta(\%)$ stands for the relative improvement after training with 20% more data in terms of average MSE over all horizons. The full result is shown in Appendix A.9.
Dataset CAPE PatchTST Dlinear MOMENT PEM
@@ 4.1 Experiment Setup
Datasets and Baselines. For pre-training CAPE, PatchTST, and PEM, we manually collected 17 distinct weekly-sampled diseases from Project Tycho (van Panhuis et al., 2018). For evaluation, we utilize five downstream datasets covering various diseases and locations: ILI USA (Centers for Disease Control and Prevention, 2023a), ILI Japan (National Institute of Infectious Diseases, 2023), COVID-19 USA (Dong et al., 2020), Measles England (Lau et al., 2020a), and Dengue across countries (OpenDengue, 2023). Additionally, RSV (Centers for Disease Control and Prevention, 2023c) and MPox (Centers for Disease Control and Prevention, 2023b) infections in the US are used to test zero-shot performance. More details can be found in Appendix A.4. We adopt baseline models from the comprehensive EpiLearn toolkit (Liu et al., 2024a), comparing CAPE against two categories: non-pretrained and pre-trained models. Non-pretrained baselines include classical statistical models such as ARIMA (Panagopoulos et al., 2021), configured in an online forecasting setting (Pham et al., 2022; Abdulmajeed et al., 2020) where parameters are updated with each new sample. We also consider deep learning models, including RNN-based approaches like LSTM and GRU (Wang et al., 2020; Natarajan et al., 2023), the linear model DLinear (Zeng et al., 2023), and transformer-based architectures (Wu et al., 2021; Zhou et al., 2021, 2022). Pre-trained baselines include state-of-the-art models such as PatchTST (Nie et al., 2022), PEM (Kamarthi & Prakash, 2023), and the time series foundation model MOMENT (Goswami et al., 2024). Further experimental details are provided in Appendix A.5.
4.2 Fine-Tuning (Full-Shot Setting)
@@
4.3 Few-Shot Performance, Zero-Shot Performance, and Ablation Study
@@
Table 3: Zero-shot performance with a lookback window length of 12. All results are averaged over 4 weeks or days in the future. $\Delta(\%)$ stands for the relative improvement over the baselines.
@@
Table 4: Ablation study.
@@
Ablation Study. We conducted an ablation study to assess CAPE’s components (Table 4). Replacing environment estimators with non-disentangling self-attention layers consistently worsened performance across all datasets, notably increasing ILI USA’s MSE from 0.309 to 0.532. Similarly, removing contrastive loss while retaining environment estimators raised Measles’ MSE from 0.269 to 0.344, with smaller increases for ILI USA and Dengue. Training CAPE directly on downstream datasets without pre-training also decreased performance, with MSE rising to 0.319 (ILI USA), 0.326 (Measles), and 1.162 (Dengue), though less than removing environment estimation. These results indicate that all CAPE components are essential for optimal forecasting.